OpenStudy (anonymous):

Combine these expressions into a single trig function or, if not possible, rewrite in terms of sin x and cos x: 1. (cos4x)(cosx)-(sin4x)(sinx) 2. sec^(4)x-tan^(4)x-2tan²x

OpenStudy (anonymous):

1st cos(5x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

cos(a+b)=cosacosb-sinasinb put a=4x nd b=x

OpenStudy (anonymous):

2nd sin^2x+cos^2x

OpenStudy (anonymous):

Can you show your process?

OpenStudy (turingtest):

do you understand the first one?

OpenStudy (turingtest):

here's what I'm using http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

Yea I got the first one now.

OpenStudy (turingtest):

so for the next we can do difference of squares on the first two terms

OpenStudy (turingtest):

\[\sec^4x-\tan^4x-2\tan^2x=(\sec^2x-\tan^2x)(\sec^2x+\tan^2x)-2\tan^2x\]

OpenStudy (turingtest):

how does this help us? look at the first set of parentheses

OpenStudy (anonymous):

It equals 1?

OpenStudy (turingtest):

yup :) so what do we have left?

OpenStudy (anonymous):

sec²x+tan²x-2tan²x

OpenStudy (turingtest):

...continue

OpenStudy (anonymous):

(1/cos²x)+(sin²x/cos²x)-(2sin²x/cos²x) (1+sin²x-2sin²x) / cos²x

OpenStudy (turingtest):

you have over-complicated things...

OpenStudy (turingtest):

sec²x+tan²x-2tan²x=sec²x-tan²x

OpenStudy (anonymous):

Then that equals 1?

OpenStudy (turingtest):

looks like it, eh? sorcery!

OpenStudy (anonymous):

sorry @enchantixPixie , i went to sleep since it was 4 in d morning here:)